Startup NIGHT 12 - 20170913 UT

Tucson: Jenny, Barry, John

Mountain: Kevin, David

Summary: Tested SX AO bin1, 2, 3, & 4 reconstructors. AGw transform verification not ready tonight. Performed Al\x92s 7 Iris D night target with the hopes that if ready tomorrow we can do the verification during the D night.

Night priorities outlined: https://docs.google.com/spreadsheets/d/1q3hQ04UESRPDJTzQn5SQ1EOXBOee_conayY7WJYgpRQ/edit#gid=0

Took some Dit5 Ndit 5 darks for Doug luci1.20170913.0032-0036 luci2.201709.0007-0011

18:33 Kevin Opening Up

Some thin clouds

18:45 test preset successful

18:48 Test Br Sky L10037 L20012

18:50 Br Gam low L10038-42 L20013-17 (5K)

18:55 H Test L10043 L2 0018

18:58 H High 11-12K L10044-48 L2 0019-23

Hlow (5-6K) L10049-53 L2 0024-28

19:04 z Test L10054 L2 29

19:06 z high L10055-0059 L2 0030-0034 Clouds moving in during these flats. No good

LUCI AO

19:13 Going to zenith, rot at 341 to Home xstage. Authorize luci1 only

19:23 Pointing check star, focus & collimating. Taking a long time to collimate

19:57 Greg lost NoMachine session to mountain

19:59. John mentioned that no high order corrections seem to be going though

20:01 Seeing 0.9\x94 on guider, We are good to go!

20:03 SX LUCI AO test - FF1725+3709 AOref 8.2mag

FATAL: Timeout slopecompcntl, Greg noted not live but is now. Resent

20:17 Illegal command for state AO Aribitrator ready. Got stuck with optimize gain. Greg kicked it and seems happy now

20:18 resending FF1725+3709. DImm 0.86\x94

Greg hit done and it safe skipped right away so 0060 is probably no good

Paused.

20:34 Greg tried to restart stuff but decided restarts of processes were needed.

\x93There is something decloaking overhead\x94 Clouds forming.

20:47 Greg took a hammer to everything. We resent preset, but pointing has drifted and we failed to reacquire guide star. David is doing a pointing check.

20:00 resending FF1725+3709 DIMM 0.88\x94

Greg has some issue with NCPA but can keep loop closed

21:12 images starting, dimm seeingg at 1.36

Aaaaand a cloud moved in after 1 exposure.

PEPSIPFU

21:18 Switching to PEPSIPFU for checking DX M3 range of adjustment for wfs pupil alignment.

21:31 preset to M5_1539 (DX AdSec ripped)

DX IE=+2 CA=+30 M3 Abs Tip = -1000 (present setting)

DX IE=+26 CA=+60 M3 Abs Tip = -1400

DX IE=+39 CA=+70 M3 Abs Tip = -1600 (negative limit of Tip for DX M3)

LUCI AO

22:24 The clouds have cleared for the most part, except some lingering in the south..so Barry a target in the south for LUCIAO to resume. reauthorize

22:25 doing pointing check and collimation 0.76\x94 on the guider, 0.9\x94 on the DIMM

22:35 Sent script for 21h31 with offaxis 11.5 Rmag off axis AO ref. Looks like we are getting some clouds. Moving to the north

22:42 Sent script to M103

22:53 W unit failed so slewed to 90, rotator to park.

22:58 going to IRAS23365+3604 (a little higher and less distance to go if W unit fails again)

23:09 running! 0.72\x94 on guider 1.1\x94 on the dimm, bin 3 L1_0062-88

Noticed that with sky offset commands that things are moving before the pause ao completes

0080 is an example

23:51 Illegal command AOArbitrator. LoopFault, IIF Offset Failed

23:53 Resumed from where script failed L1 0090-0095 Seeing 0.73 on guider, 0.88 on Dimm

COMPLETED Bin 3 target

0:07 Slew to IC 5117 AOref 9th mag, Bin 2 (right on the border of 1,2) 153 modes

L1 0096-0122

Observer gui stopped updating on obs2 and rm525-1 at the same time (~0:19). [observer gui

stopped updating previous night, too] Took luci snapshot on both. Exposures continued.

1:00 Failed. Continued but with edited script (6\x94 offsets instead) = failed, had to resend preset.

1:04 Observer panels froze on obs2 and rm525-1. Resent to IC 5117. Bin 1 with NCPA actually working. L1 123

Loop opened after during 0127 because of seeing bump

Closed and restarted (0.52\x94 guider, 0.79\x94 on DIMM)

1:52 Safe skip during offset. Last image 00145

1:54 resent preset, error: left secondary hexapod not in position

2:07 Resent preset Bin1 L1 0146, skipped to J where script died.

2:10 AO arbitrator process croaked. L1 0148 loop open. Moving on (seeing 0.64\x94 guider)

2:13 Going to 7 Iris, non sidereal ACE track preset. Preset failed, slopecompcntl timeout

2:17 resent to 7 Iris, DIMM 0.91\x94, Blasted with BrGam. Switching to H2

2:33 Sent to 7 Iris H2 setup, Bin1, 400 modes, NCPA on 2.51x20 L1 0154-0163

2:46 FF0234 AOref of 14.5 mag, testing bin4, 36 modes Seeing 0.85\x94 on DIMM,

Start L1 0164

0166 probe off axis guider moved just out of parol field (who cares, maybe save for testing software)

0171 also probe off axis guider moved just out of parol field (offset abs x -7 did it I think)

0.87\x94 Dimm, 0.61\x94 on guider

03:16 - 0.89\x94 DIMM 0.64\x94 guider - airmass 1.2

0180 also probe off axis guider moved just out of parol field

Last image 0181

3:22 7 Iris again with H2 setup L1 0182-0185 DIMM 0.81\x94 Seeing

Not sure if L1 0186 is ok Offset XY command interrupted

03:43 - Asteroid was on edge of patrol field - suggests pointing is off

03:45 - John H. manually moved to get the target closer - Greg running RunAO so we can continue script without resending preset

3:52 resumed 7 Isis from where we left off. L1 0187-192

4:15 Standard SJ9105 H & H2 L2 0193-202

0196

Binary possibly, Surprise! On edge of correction so more observation needed.

4:27 7 Iris, one last go. Jump in 203 L1 0203-212


Closed up and handed over to Dave for LUCI1 Wavelength calibration

Comments

 
Topic revision: r2 - 12 Oct 2017, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback